OpenStudy (anonymous):

Complete the subtraction problem below. - = ? A. 2 B. 1/50 C. 2/5 D. 3/50

OpenStudy (anonymous):

wait this was the question

OpenStudy (anonymous):

4/5 - 1/5= ?

OpenStudy (anonymous):

and those are the options

OpenStudy (anonymous):

i think its C*

OpenStudy (anonymous):

does that say 3/50 in D?? or 3/5?

OpenStudy (agent0smith):

^^^ yeah there's no correct answer in the choices

OpenStudy (anonymous):

3/50

OpenStudy (agent0smith):

Well \[\Large \frac{ 4 }{ 5 }-\frac{ 1 }{ 5} = \frac{ 4-1 }{ 5 }\]

OpenStudy (agent0smith):

I like that 2 of the answer choices aren't even remotely correct, and how would anyone ever even decide on one of those answers...